write the equation of the line passing through each of the given points with the indicated slope. (rewrite in standard from) Give your results in slope-intercept form, where possible.
(-1,3) and (4, -2)

Answer by checkley75(3666)

SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(-2-3)/(4+1)
SLOPE=-5/5=-1
NOW REPLACE THE X & Y TERMS WITH -1 & 3 & SOLVE FOR THEY INTERCEPT b
Y=mX+b
3=-1*-1+b
3=1+b
b=3-1
b=2 THE Y INTERCEPT.
Y=-X+2 IS THE LINE EQUATION.